### v1

# Automated Detection of Galactic Rings from SDSS Images

Abstract
Morphological features in galaxies, like spiral arms, bars, rings, tidal tails etc. carry information about their structure, origin and evolution. It is therefore important to catalogue and study such features and to correlate them with other basic galaxy properties the environment in which the galaxies are located and their interactions with other galaxies. Surveys such as SDSS, Pan-STARRS, HSC-SSP have made available very large samples of galaxies for gainful morphological studies. The availability of galaxy images and catalogues will increase manifold with future surveys like LSST. The volume of present and future data is so large that traditional methods, which involve expert astronomers identifying morphological features through visual inspection, are no longer sufficient. It is therefore necessary to use AI based techniques like machine learning and deep learning for finding morphological structures quickly and efficiently. We report in this study the application of deep learning for finding ring like structures in galaxy images from the Sloan Digital Sky Survey (SDSS) data release DR18. We use a catalogue by Buta (2017) of ringed galaxies from the SDSS to train the network reaching good accuracy and recall, and generate a catalogue of 29420 galaxies of which 9805 have ring like structures with prediction confidence exceeding 90 percent. Using a catalogue of barred galaxy images identified by Abraham et. al. (2018) using deep learning techniques, we identify a set of 2087 galaxies with bars as well as rings. The catalogues should be very useful in understanding the origin of these important morphological structures.
Astronomy data analysis (1858), Astronomy image processing (2306), Catalogs (205), Galaxies (573)
1 Introduction
Galaxies display a wide variety of forms as a consequence of the differences in their intrinsic structure, interaction with other galaxies as well as observational biases. Efforts to make sense of galaxy morphology with its multitude of forms have a long history dating back to Edwin Hubble and others. The Hubble classification system (Hubble, 1926) arranged galaxies in a sequence ranging from ellipticals to lenticulars to spirals to irregulars shaped objects. Galaxies other than the ellipticals were further classified into those with bar like structures and those without such structures, resulting in a tuning fork like diagram. Other more detailed classification systems have been developed (see, e.g., Van den Bergh, 1998), taking into account further morphological features.
In recent decades, the availability of many millions of higher resolution multiwavelength images from large deep field surveys made using modern instrumentation have made possible detailed morphological studies (Buta, 2011). Galaxy morphology plays a key role in bettering our understanding of the secular processes that underlie galaxy evolution. Several important questions concerning the formation and evolution of galaxies can be addressed from the observed change in morphology with epoch.
A fairly common feature of disk galaxies is the presence of a ring shaped pattern in their light distribution (Buta & Combes, 1996). These rings in galaxies have come to be considered an integral part of galaxy morphology (Kormendy, 1979). Recognizing their importance, the de Vaucouleurs Revised Hubble-Sandage Classification System (VRHS; De Vaucouleurs, 1959) added rings as another dimension to the two-dimensional Hubble tuning fork, turning it into a classification volume. Rings in galaxies can fundamentally be divided into normal rings, also called ‘resonance rings’ and ‘catastrophic rings’, which are a result of galaxy collisions (Buta & Combes, 1996). Although different subclasses exist within the catastrophic rings, such as polar, accretion, and collisional rings, they constitute a very small minority of all observed rings. Madore et al. (2009) have estimated the abundance of collisional rings to be only 1 in 1000. Several theories have been put forward to explain the different ring structures that are observed, with the resonance interpretation being the most popular. The “Manifold Theory” that has been proposed recently Athanassoula et al. (2009) also has had considerable success in explaining several aspects of ring morphology.
Several catalogues that contain the visual morphology classification of galaxies are available today. Notable among them are Nair & Abraham (2010) and Fukugita et al. (2007), with the former containing the detailed classifications of 14,034 galaxies. Crowd-sourcing efforts such as the Galaxy Zoo citizen science project (Lintott et al., 2011) have managed to ramp up the number of galaxy morphology classifications to nearly 900,000 by involving volunteers. This vast database has motivated others to pick specific subtypes of galaxies for detailed classifications. For example, Buta (2017) has exhaustively classified 3962 ringed galaxies taken from this set. However, all these efforts involve a human in the loop for classifying galaxy morphologies.
With an explosion in the number of galaxy images being produced as part of imaging surveys like the Sloan Digital Sky Survey (SDSS; York et al., 2000), Panoramic Survey Telescope and Rapid Response System (Pan-STARRS; Kaiser et al., 2002), Dark Energy Survey (DES; Dark Energy Survey Collaboration et al., 2016) etc., automated methods of morphology classification have become very important. Although traditional image processing techniques, many of which are histogram-based, have had some success (see, e.g., Shamir, 2020), these are inappropriate for complex tasks such as shape recognition from images. However, the advent of artificial intelligence (AI) techniques like Convolutional Neural Networks (CNNs; LeCun et al., 1995) has significantly improved the automated efforts for galaxy morphology classification. One of the initial efforts in this area have been made by Banerji et al. (2010). Dieleman et al. (2015) have used CNNs to create a model that can match the combined consensus of the Galazy Zoo volunteers on all of the questions regarding galaxy morphology that are part of the survey. Abraham et al. (2018); Sánchez et al. (2018) have used CNNs for the automated classification of galaxy morphological features such as bars, bulges, edge-on morphology etc. More recently, Shimakawa et al. (2023) have used a deep-learning model to obtain ring classification for a very large sample of galaxies taken from the Hyper Suprime-Cam Subaru Strategic Program.
In the present study we investigate the efficacy of an automated method based on Convolutional Neural Networks for the detection of rings in galaxies. We use galaxy images from the Sloan Digital Sky Survey (SDSS; York et al., 2000) for the training and testing of our network. We treat the problem as a binary classification one, treating all sub-classes of rings as a single ring type and distinguishing it from galaxies that do not contain any rings, the non-ring type. The trained network can be used to identify ringed galaxies from the SDSS.
The rest of the paper is organized as follows: In Section 2, we describe the data that we used. The neural network architecture is introduced in Section 3. Sections 4 and 5 describe the data augmentation techniques and the training procedure for our neural network respectively. Section 6 analyses our results on the trained network. Section 7 presents the catalog of ringed galaxies generated using our trained network. Section 8 presents a subset of the ringed galaxies that have also been identified to have bars. Finally Section 9 gives the concluding remarks.
2 Data
The performance of a supervised machine learning classifier depends to a large extent on the quality of the labelled data that is made available to it. For deep learning models the quantity becomes important as well because of the large number of trainable or free parameters in the model. Many of the standard datasets available for galaxy morphology classification do not contains rings as a class (see, e.g., DECALS10  [^1] https://astro.utoronto.ca/~hleung/shared/Galaxy10/Galaxy10_DECals.h5). Even with the Galaxy Zoo 2 dataset (Willett et al., 2013), ringed galaxies for which at least 50 percent of zoo volunteers agreed on a response are relatively small in number.
2.1 Rings and Non-Rings
We used the Buta (2017) catalog as our primary source to identify galaxies with rings. The author has provided detailed visual morphological classification of 3962 galaxies with rings. The classifications are done within the framework of the Comprehensive de Vaucouleurs revised Hubble-Sandage system (CVRHS; Buta et al., 2015, 2007). The catalog also provides the author’s comments that are useful for creating a good quality training set by identifying rare cases of rings, poorly resolved galaxies etc. The galaxies used for classification by the author were picked from the Galaxy Zoo 2 (Willett et al., 2013) survey (hereafter, GZ2).
Along with a good quality training sample for rings, it is equally important to obtain a good sample of galaxies that do not have any presence of rings. Since the Buta (2017) catalog does not contain galaxies without rings, we used the Nair & Abraham (2010) catalog, which is one of the largest catalogs for the visual morphological classification of galaxies. They used galaxy images from the SDSS DR4 release (Stoughton et al. (2002); Adelman-McCarthy et al. (2006)). Galaxies with spectroscopic redshifts in the range $0.10.01<z<$ 0.1 and extinction corrected g’-band magnitude $<16$ were selected from the spectroscopic main sample (Strauss et al., 2002). This led to their final sample size of 14,038 galaxies. Figure 1 shows the actual distribution of redshift and extinction corrected g-magnitude for the ringed and non-ringed galaxies in our training set.
Figure 1: The normalized distribution of redshifts and extinction corrected g-mag for the ringed galaxies(left) and non-ringed galaxies(right)
The classification scheme used in Nair & Abraham (2010) is not the CVRHS. Rather it is primarily based on the Carnegie Atlas of Galaxies (Sandage & Bedke, 1994) used in consultation with the Third Reference Catalog of Bright Galaxies (RC3; de Vaucouleurs et al., 1991). In order to use only galaxies without rings, we selected from the catalog galaxies that have the ring type and ring flag column set to zero. Along with ringed galaxies identified by Buta (2017), we also used the ringed galaxies identified by Nair & Abraham (2010). Their catalog has identified 2598 rings from a total of 14034 galaxies. By visual inspection we identified 800 galaxies that had clearly identifiable rings and added this to our training sample.
2.2 Selection Criteria
The images used for the GZ2 survey come from the SDSS Data Release 7 (DR7; Abazajian et al., 2009). The GZ2 includes galaxies with extinction corrected Petrosian half-light magnitude in the r-band < $17.0$, Petrosian radius $\tt{petroR90_{r}}>3$ arcsec and spectroscopic redshift in the range $0.250.0005<z<$ 0.25, when it is known. Finally, galaxies that have SDSS flags which are either SATURATED, BRIGHT, OR BLENDED without an accompanying NODEBLEND were removed. We further selected galaxies that have an extinction corrected g’-band magnitude < 16 based on the selection criteria used for the non-ring galaxies. The spectroscopic redshift was also limited to be in the same range as that of the non-ring galaxies, specifically $0.10.01<z<$ 0.1.
2.3 SDSS Image Cutouts
Most imaging surveys provide an Application Programming Interface (API) access for downloading image cutouts in bulk from their servers. For our purposes we needed to download images in bulk both for creating the training data as well for catalog preparation. We modified a python library [^2] https://pypi.org/project/panstamps/ that already exists in the public domain for downloading images. We incorporated multi-threaded parallel downloading using the Concurrent  [^3] https://docs.python.org/3/library/concurrent.futures.html python module into the script. This enabled us to download images in parallel, speeding up the whole data downloading process. The script takes as input a comma-separated-value (CSV) file containing a minimum of three columns, the Right Ascension (RA), Declination (DEC) and a unique key that can be used to name the downloaded images. Figure 2 shows a sample of the galaxies with and without rings in our training set.
Refer to caption Galaxies with rings
Refer to caption Galaxies with no rings
Figure 2: A sample of galaxies with and without rings from our training set.
3 Network Design
AlexNet (Krizhevsky et al., 2012) is one of the earliest CNN architectures desgined and yet one of the most reliable. Having a smaller size than most modern architectures results in lesser training time and also the requirement of lesser amount of training data. The original AlexNet architecture was designed for the ImageNet challenge (Deng et al., 2009). The network has eight layers with the first five being convolutional layers and the remaining three being fully-connected. The first convolutional layer filters the 224 × 224 × 3 input image with 96 kernels of size 11 × 11 × 3 with a stride of four pixels. All of the convolution and fully connected layers are immediately followed by a block consisting of a Rectified Linear Unit activation (ReLU; Nair & Hinton 2010) and a batch normalisation (NORM) layer. The first, second and final convolution layers are each followed by a maxpooling layer, that leads to 50 percent downsampling of the convolution feature maps. These layers also contribute to the rotational invariance of the learnt model weights (Boureau et al., 2010). The max-pooling layers of the first two convolutions are followed by dropout layers (Srivastava et al., 2014) having a dropout fraction of 0.25. These prevent overfitting of the model by randomly dropping 25 percent of the weights during each training iteration.
We left the input layer size unchanged since it was suitable for our purposes. However, galactic rings are features that are harder to distinguish in comparison to the classes in the ImageNet challenge. Hence we used a smaller size of five for the kernels in the first convolutional layer. This seemed to have a considerable impact on the accuracy of the network. Thus in our network the first convolution layer has 96 kernels of size $5\times 5\times 3$ and the output is fed to the second layer of convolution with 256 kernels each of size $5\times 5\times 48$ The third, fourth and fifth convolutional layers have respectively 384, 384 and 256 kernels of size $3\times 3\times 256$, $3\times 3\times 192$ and $3\times 3\times 192$ respectively. These learn more complex features whose final output is fed into another max-pooling layer. The output of the final pooling layer is then fed into a series of fully connected layers of size 4096. These layers are connected by dropout layers which have 0.50 dropout fraction. The final layer is the sigmoid activation function layer which outputs a value that can be interpreted as the probability of the input belonging to the positive class. The exact architecture of our network is shown as a table and schematic diagram in Figure 3.
Name Type Shape
0 conv2d Conv2D (None, 100, 100, 96)
1 activation Activation (None, 100, 100, 96)
2 batch_normalization BatchNormalization (None, 100, 100, 96)
3 max_pooling2d MaxPooling2D (None, 49, 49, 96)
4 dropout Dropout (None, 49, 49, 96)
5 conv2d_1 Conv2D (None, 49, 49, 256)
6 activation_1 Activation (None, 49, 49, 256)
7 batch_normalization_1 BatchNormalization (None, 49, 49, 256)
8 max_pooling2d_1 MaxPooling2D (None, 24, 24, 256)
9 dropout_1 Dropout (None, 24, 24, 256)
10 conv2d_2 Conv2D (None, 24, 24, 384)
11 activation_2 Activation (None, 24, 24, 384)
12 batch_normalization_2 BatchNormalization (None, 24, 24, 384)
13 conv2d_3 Conv2D (None, 24, 24, 384)
14 activation_3 Activation (None, 24, 24, 384)
15 batch_normalization_3 BatchNormalization (None, 24, 24, 384)
16 conv2d_4 Conv2D (None, 24, 24, 256)
17 activation_4 Activation (None, 24, 24, 256)
18 batch_normalization_4 BatchNormalization (None, 24, 24, 256)
19 max_pooling2d_2 MaxPooling2D (None, 11, 11, 256)
20 dropout_2 Dropout (None, 11, 11, 256)
21 flatten Flatten (None, 30976)
22 dense Dense (None, 4096)
23 activation_5 Activation (None, 4096)
24 batch_normalization_5 BatchNormalization (None, 4096)
25 dropout_3 Dropout (None, 4096)
26 dense_1 Dense (None, 4096)
27 activation_6 Activation (None, 4096)
28 batch_normalization_6 BatchNormalization (None, 4096)
29 dropout_4 Dropout (None, 4096)
30 dense_2 Dense (None, 2)
31 activation_7 Activation (None, 2)
Figure 3: A tabular description of the AlexNet model architecture that we have used, generated with the Keras package (above) and a visual representation of the same (below). Code for generating the figure was adapated from Iqbal (2018)
3.1 Selection of Hyperparameters
Hyperparameter tuning is an important part of any machine learning or deep learning project. Hyperparameters are those parameters that are not learnt or updated during the learning process and hence need to be manually set by the programmer. These can include the network architecture itself and several other parameters like the batch size used for the stochastic gradient descent or the number of epochs of training etc. These parameters have a direct bearing on the performance and accuracy of the model. Finding the right set of hyperparameters is usually a trial and error process.
Since a machine learning model learns iteratively by making errors the loss function is an important hyperparameter to tune. The loss function specifies what constitutes an error and how much weightage should be given to individual errors. We have used categorical crossentropy [^4] https://www.tensorflow.org/api_docs/python/tf/keras/metrics/categorical_crossentropy, which is a standard loss function suitable for binary classification problems. The optimizer function has an effect on how the network weights are updated in order to minimize this loss. We used the Adam optimizer with a learning rate of 0.003. Regularization techniques are a set of safeguard measures that prevent the model from overfitting on the training data and thus losing its ability to generalize. We have used L2 weight regularization with a factor of 0.0002 to penalize the weights from growing too much during training. The batch size controls how many input samples are considered together for evaluating the loss function. This leads to a trade-off between accuracy and speed. A low batch size leads to better accuracy often at the cost of very large training times. A batch size of 16 was seen to be sufficient in our case. The number of epochs or (num epochs) decides how many times the network gets to see the complete training data. Early Stopping (Prechelt, 1998) is a technique that can be used to stop the training process if subsquent trainings do not improve the metrics we are interested in like the validation loss. However, using it can lead to uncertaininity as it introduces another hyperparameter called the patience that needs to be properly set. We trained the network for num epochs without early stopping.
4 Image augmentation
Neural networks are prone to overfitting if the data available for training is insufficient. The size of data required increases with the complexity of the network which in turn is dependent on the number of free parameters available for tuning. Increasing the training data is limited by the availability of dedicated catalogs that can provide a sufficiently large number of ringed galaxies. A standard practice in such cases is to create more examples for training the network through data augmentation. This involves the transformation of existing images to generate realistic looking new images. CNNs are translation invariant by design but they are not rotation invariant. Additionally, individual CNN filters although translational invariant are not invariant to rotations of the images. However ringed galaxies have rotational symmetry and hence their classification should be invariant to rotations. By transforming each of the training images using random rotations we could increase the dataset size at the same time making the network invariant to rotations. We have used operations such as horizontal and vertical flipping, rotation through arbitrary angles, brightness and contrast adjustments for augmenting our training set. The augmentations that we have used are illustrated in Figure 4 using a random galaxy.
We implemented image augmentations using the on-the-fly mode provided by the keras-tensorflow (Chollet et al., 2015) framework . In this way, during training, each image from the training set is taken and a predefined sequence of transformations are applied to it. The parameters used for the transformations are randomized so that each epoch, the network sees a different set of augmented images. In this way we avoid having to generate all the augmented images before the start of the training and also save a lot of disk space. Additionally by making use of such a data pipeline we can make the process of k-fold splitting of the dataset easier to manage.
Refer to caption Original
Refer to caption Horizontal Flip
Refer to caption Vertical Flip
Refer to caption Rotation
Refer to caption Brightness
Refer to caption Contrast
Figure 4: The different types of image augmentations that are done to each image in our training set.
5 Training
The galaxy image cutouts used for training are 3 channel RGB images with dimensions of $256\times 256$ pixels. The images are dynamically resized by the Keras image_dataset_from_directory  [^5] https://www.tensorflow.org/api_docs/python/tf/keras/utils/image_dataset_from_directory function into the required input shape of $240\times 240$ before being fed into the network. The pixels intensities in the range (0-255) are also rescaled to the range (0,1) for easier computation. The input size and other parameters for the network like num epochs, batch size are controlled through a config file which is read at run time. The images are first resized to the desired input size and then loaded into memory in nbatches using the Keras image_dataset_from_directory function. A generator avoids the need for all of the images to be loaded into memory at once.
5.1 K-fold Cross-Validation Training
We decided to do the training using k-fold cross-validation (hereafter, CV) split with $k=5$. This is better suited for evaluating our model performance since our dataset size is limited. The k-fold CV technique makes sure that we are using the entire available data for testing our model. This is achieved by first splitting our dataset into k different parts or folds and running k different training sessions. Each model is trained using k-1 folds of the data and tested on a single fold. The evaluation metrics computed on the single fold are then averaged over the k different training sessions. The best performing model can then be chosen for further analysis. Note that when using the k-fold CV technique there is a significant trade-off in terms of total computation time.
The scikit-learn (Pedregosa et al., 2011) python library has a train-test split method that can be used to randomly shuffle and split a given array into two sets by accepting a split fraction parameter. However using this was impractical as it required the whole dataset to be loaded into memory at once. Besides using on-the-fly data augmentation meant that the data is generated dynamically during training and was not available before hand. To overcome both these issues we used the scikit-learn library to first split the data into k partitions on disk. We then used various combinations of these partitions to create the k datasets required for k-fold CV training. These datasets could now be used with the Keras data loading generators for batch-wise loading of the data and selective data augmentation could be done on the training data alone.
The network was trained for a total of 140 epochs. The training graph in Figure 5 shows the variation of the categorical_crossentropy loss function with increasing epochs for both the training and validation data. It gives provides insight into the learning process happening within the network. Among other things, it is useful in identifying whether the network is overfitting on the training data. When this happens the validation loss of the network increases whereas the training loss keeps decreasing. In our case, the training loss and validation loss are seen to decrease together with each epoch and there is no significant overfitting that occurs. There is further scope for training the network albeit with minimal returns.
Class Precision(%) Recall(%) F1-Score(%)
Rings 93 89 91
Non-Rings 99 99 99
Figure 5: The training graph, which shows a plot of the training and validation loss against the number of epochs (top-left). The confusion matrix, which shows a graphical representation of the number of true positives, true negatives, false positives and false negatives in our test set (top-right). The classification report showing the precision and recall for both rings and non-rings in our test set (bottom).
The training was done on the Amazon Web Services platform on an Intel Xeon workstation with 4 virtual CPUs and 16 GB RAM making use of a Tesla (g4dn.xlarge) GPU accelerator with 16 GB graphics memory. A community AMI (Amazon Machine Image) running Arch Linux was modified to function as our base operating system. The data and model are version controlled using Data Version Control (DVC; Kuprieiev et al., 2022) and stored on Amazon S3 cloud storage. The code itself is version controlled and made available on github  [^6] https://github.com/linnabraham/galactic-rings.git. The entire training session takes 2.5 hours to run with our hardware specifications.
6 Results
6.1 Evaluation Metrics for Unbalanced Datasets
Numerous metrics are available in the literature for evaluating the performance of a neural network. However choosing the right metrics to evaluate depends on the specific task at hand. Often several metrics have to be considered together to get the complete picture. Accuracy is a common metric used in binary classification problems. It is the number of correct classifications over the total number of classifications made, expressed as a percentage. Mathematically, it can be expressed as
$\rm{accuracy}(f)=\frac{\Sigma_{x_{i}\in X_{test}}1\{f(x_{i})=y_{i}\}}{X_{test}}$
where $f(x_{i})$ is the label assigned to an image from our test set by our model $f$ and $y_{i}$ is the true label associated with that image. However, accuracy is a class agnostic measure because it treats both classes on an equal footing. Hence depending on the degree of class imbalance in the test set, having a high enough accuracy might not be sufficient to consider a classifier as performing well. In such cases, it might be helpful to also consider additional metrics that complement accuracy. Precision and recall are two such metrics that can be used for this purpose and are also class specific.
Precision and recall are defined as
$\rm{precision}=\frac{TP}{TP+FP}$
$\rm{recall}=\frac{TP}{TP+FN}$
Here TP denotes the number of true positives, FP the number of false positives and FN the number of false negatives. The precision and recall measure how well the classifier is performing with regard to a particular class. Our trained model obtained a classification accuracy of 98 percent on the testing set. The classification report shown in Figure 5 summarizes the precision and recall we have obtained for both the classes. The precision obtained for the rings was 93 percent. Our model also has a good recall value of 89 percent.
It should also be noted that the precision and recall should not be viewed in isolation. It is trivial to obtain perfect scores for either precision or recall by predicting all examples as belonging to the positive class. Such a classifier can be called a no-skill classifier. Also for a particular model there is often a trade-off between the precision and recall. Note that the two of the metrics that are interesting for the astronomer in relation to astronomical catalogs are its purity and completeness. The former is related to the precision of the model and the latter to the recall. Using the predicted probabilities of our network, a catalog with higher purity can be obtained by appropriately selecting the classification threshold. However it is to be kept in mind that higher purity comes at a cost of reduced completeness. The nature of the trade-off can be understood using the precision-recall curve shown in Figure 6.
The Receiver Operating Characteristics curve or the ROC curve can be thought of as a way to compare the performance of different classifiers. Each classifier contributes a single point to the plot. However even for a single classifier we can use different thresholds to get a curve. The ROC curve shown in Figure 6 plots the True Positive Rate (TPR) against the False Positive Rate (FPR) at different thresholds, where the TPR and FPR are given by the following equations:
$\rm{TPR}=\frac{TP}{TP+FN}$
$\rm{FPR}=\frac{FP}{FP+TN}$
Figure 7 shows galaxies in the validation set that are correctly classified (True Positives and True Negatives) and failures (False Positives and False Negatives) obtained using our trained network.
Figure 6: On the left, the Receiver Operator Characteristics (ROC) curve which shows the plot of the False Positive Rate (FPR) against the True Positive Rate (TPR) obtained by varying the classification threshold. On the right, the Precision-Recall curve which shows the plot of precision against recall for ringed galaxies in our test set computed at various thresholds. In both cases, the dashed line shows the curve expected from a no-skill classifier.
Refer to caption True positives from our validation set
Refer to caption True negatives from our validation set
Refer to caption False positives from our validation set
Refer to caption False negatives from our validation set
Figure 7: Examples of images from our validation set that are both correctly and incorrectly classified.
6.2 Prediction at Different Zoom Levels
To get a better insight into what the model has learnt we downloaded a single galaxy at different zoom levels using the SDSS image cutout service. This is done by varying the scale factor and downloading the different images. The trained network was then used to predict on these images. Figure 8 shows different versions of a particular ringed galaxy along with the probability score returned by the network. It can be seen that probability score goes to zero both when the images is too zoomed out or zoomed in such that the ring feature is not visible. For other zoom levels the probability score given by the network are comparable to what a human observer would assign.
Figure 8: Cutouts of a single galaxy downloaded at various zoom levels from the SDSS cutouts server and predicted using the trained network. The prediction scores obtained for each are shown in red color.
6.3 Analysis of Failed Predictions
To get a better idea of the performance of our classifier we analyzed the failed predictions. A quantitative way of doing this is to compute the log-loss for each sample in our test set. In a binary classification problem such as ours, with a sigmoid activation for the last layer, the predicted value is a real valued number between 0 and 1. Hence it is often interpreted as a probability of the sample actually belonging to the positive class. However for the negative class a low predicted probability implies a high confidence value. Hence in this case we might need to take the difference of the value from one in order to get the real confidence. When analyzing the failures in our testing set we can use a loss function that is weighted using the confidence with which a prediction has been made. The log-loss is a particular kind of loss function that is suitable for this purpose.
$\rm{logloss_{i}}=-[y_{i}\ln p_{i}+(1-y_{i})\ln(1-p_{i})]$
Here $y_{i}$ is the true value and $p_{i}$ is the predicted probability score for observation $i$. The log-loss can be computed for each sample in our test set. The histogram of these losses can then be used to get a better idea about the overall performance of the model. Figure 9 shows the histogram of log-loss obtained for our test set. The peak of the histograms are towards the lower end. Which means that there are less number of wrong predictions made with a high confidence value and more number of correct predictions with a high confidence value. That is, it can be seen that the confidence of wrong prediction is much lower than that of the correct predictions. The figure also shows representative images images from our test set corresponding to the False Negatives and having a high log-loss values. The rings in these galaxies are not very distinct and one of the galaxies have a high inclination angle that might have confused the network.
6.4 Model Interpretability
While the model is being trained, it tries to learn an aggregate representation for the whole data. Once the model is trained, if we use it to make predictions on the training data itself, we will be able to get some idea of what the model has learnt. Such techniques come under the domain of model interpretability. We used the trained model to predict on the ringed galaxies in our training set. Since rings belong to the positive class we expect a perfect prediction of one, but which is not usually the case. Hence we compute the difference from one of the actual prediction as our error. This error can be used for finding subsets of ringed galaxies in our training sample that are either easy for the network to learn or difficult to learn. The results are shown in Figure 10. It can be seen that among the galaxy images that the network had trouble in learning there are many with image artefacts or the presence of background stars or galaxies. Many of the galaxies have a zoomed in view as well. The galaxy images that were easily identified by the network had a better contrast with the background and the galaxies were at a normal zoom level.
Figure 9: The normalized distribution of log-loss in our test set (above). The true negatives are shown separately because of the difference in bin range. A sample of galaxies from the false negatives having the highest log-loss values (below).
Refer to caption Galaxies with a low loss
Refer to caption Galaxies with high loss
Figure 10: Ringed galaxies from our training sample that have a low loss (above ) and high loss (below)
7 A Catalog of Ringed Galaxies
To prepare a catalog of ringed galaxies from SDSS we used the same selection criteria that was used for the training data preparation. The SDSS data can be queried using the CASJOBS  [^7] https://skyserver.sdss.org/casjobs/default.aspx interface. The complete query is given in the Appendix A. The source list obtained from the CASJOBS server was downloaded using our bulk download script. The trained network was then used to predict on these galaxies. A preview of the catalog is included in Table 1. Figure 13 shows a selection of galaxies with rings taken from our catalog.
In Figure 12 we have shown the histogram of the prediction scores or confidence values for all the galaxies in the prediction set. The sharp peak in the histogram near zero arises because there are far more non-rings than rings. In order to do a more thorough testing of our network on real world data, we show in Figure 11 an image mosaic created from 50 randomly selected galaxies from our prediction set along with the machine generated probability scores. The galaxies are also sorted in descending order of the prediction score. Based on the natural abundance of ringed galaxies we may expect less than 10 of these to be ringed. It can be seen on visual inspection that most of the galaxies with true rings have a high prediction scores in comparison to the non-rings and hence they have clustered towards the top of the image mosaic. We found 9805 galaxies that are predicted to have rings using a classification threshold of 0.90.
Figure 11: A mosaic of 50 randomly selected images from the prediction set sorted using the predicted probability of having a ring. Visual inspection shows that the rings have accumulated towards the bottom and the non-rings towards the top.
Figure 12: The histogram of probabilities from our ring catalog. Note that the y-axis uses a log scale. The plot shows a sharp peak near 0 corresponding to galaxies which have been predicted as non-rings with very high confidence. And another peak near 0.90 corresponding to galaxies which were predicted as rings with high confidence.
ra dec objid gmag deVRad_r deVAB_g redshift Prediction Label
209.103777 -0.247728 1237648704051478528 14.155170 10.535850 0.836304 0.029755 0.989081 Rings
209.114487 -0.287759 1237648704051478784 15.777170 25.383930 0.343417 0.029818 0.000178 NonRings
209.529402 -0.392480 1237648704051675136 14.630820 15.223800 0.894546 0.029576 0.562100 NonRings
208.686043 0.115120 1237648704588153088 15.209120 16.307050 0.669667 0.029839 0.915272 Rings
218.916989 0.204939 1237648704592609280 15.854450 3.193142 0.782952 0.029226 0.947405 Rings
221.078715 0.041915 1237648704593592320 15.101800 4.807141 0.660744 0.028935 0.947263 Rings
221.076526 0.044058 1237648704593592320 15.323840 21.666000 0.570289 0.028187 0.010766 NonRings
218.832171 0.475730 1237648705129480192 15.486240 4.657629 0.618569 0.029567 0.631004 Rings
221.263726 0.526979 1237648705130529024 15.389920 29.667630 0.462591 0.028732 0.073611 NonRings
199.295541 0.913565 1237648705657765888 15.475490 15.956910 0.589317 0.080584 0.496464 NonRings
217.502392 1.046087 1237648705665761280 15.427110 4.664006 0.631887 0.029724 0.664599 Rings
221.954869 1.051587 1237648705667727616 15.943330 14.861050 0.949775 0.028619 0.000456 NonRings
232.722701 0.994218 1237648705672446208 15.644950 11.673120 0.350001 0.029820 0.193692 NonRings
134.447161 -0.199955 1237648721210769664 14.868640 7.398778 0.888698 0.028206 0.978550 Rings
157.940531 0.387938 1237648721757929472 15.397520 4.335503 0.967293 0.028369 0.895102 Rings
159.239044 0.229754 1237648721758519552 14.453670 19.212620 0.778882 0.029224 0.954128 Rings
220.742692 0.268993 1237648721785454592 15.710900 2.857519 0.672701 0.028958 0.806524 Rings
199.509777 1.114085 1237648722849890560 15.857810 5.050115 0.768515 0.080753 0.590488 NonRings
176.371717 -3.518446 1237650369406631936 15.433640 1.949464 0.795187 0.028688 0.782616 Rings
177.410003 -3.508220 1237650369407090688 15.904690 14.294680 0.682261 0.028744 0.231090 NonRings
ra dec SDSS_Objid Prediction Label
$0$1.113000 1237663275780276480 0.001809 Barred-NonRings
0.019800 0.781700 1237657191978959104 0.650422 Barred-Rings
$0$0.723700 1237663783123681280 0.039325 Barred-NonRings
$0$1.213600 1237663275780276480 0.060768 Barred-NonRings
0.114500 14.962500 1237656495650570496 0.208050 Barred-NonRings
0.117000 14.381200 1237652942638481664 0.183115 Barred-NonRings
0.128700 -1.213000 1237663275780341760 0.532812 Barred-Rings
0.158900 14.623600 1237656495113699328 0.024679 Barred-NonRings
0.166500 -1.191200 1237663275780342016 0.036194 Barred-NonRings
0.228100 15.218700 1237652943712289024 0.000647 Barred-NonRings
0.241000 14.190400 1237656494576894208 0.162884 Barred-NonRings
0.241700 14.864400 1237652943175418112 0.215693 Barred-NonRings
0.261300 -0.257100 1237663783660617728 0.001189 Barred-NonRings
0.297800 15.985900 1237656496724443392 0.000518 Barred-NonRings
0.300100 15.351400 1237652943712289024 0.173043 Barred-NonRings
$0$0.992900 1237657189831606272 0.008202 Barred-NonRings
0.368800 -0.224400 1237663783660683520 0.321358 Barred-NonRings
0.380900 14.407500 1237652942638612736 0.211441 Barred-NonRings
0.388600 -0.730500 1237663783123812608 0.000156 Barred-NonRings
0.398900 -0.395200 1237663783660683520 0.303563 Barred-NonRings
Table 1: A preview of the catalog of ringed galaxies (above) and of barred rings (below) with the first 20 rows from each.
8 A Catalog of Ringed galaxies with Bars
Bars in galaxies are much more common than rings. Approximately half of the observed galaxies are seen to have bars Knapen (1999). However galaxies that contain both bars and rings are certainly a smaller fraction. Dedicated catalogs for such galaxies are also quite rare. Here we use the Abraham et al. (2018) catalog, which is an automated catalog of galaxies with bars in SDSS, to find galaxies with both bars and rings. The Abraham et al. (2018) catalog contains 111,838 galaxies that have been predicted as either barred or unbarred. We selected the barred galaxies which came to be 25,781 in number and used our trained network to predict which of these galaxies also have rings. Figure 13 shows a random selection of the barred galaxies that have been predicted to have rings with a probability score of greater than 90 percent. We found 2087 galaxies that contained both bars and rings using a classification threshold of $0.5$. A preview of the catalog containing barred galaxies with predictions for the presence or absence of a ring is shown in Table 1.
Refer to caption Galaxies with rings
Refer to caption Galaxies with bars and rings
Figure 13: Some good examples of ringed galaxies (above) and barred ring galaxies(below) from our catalog.
9 Conclusion
We have shown that a relatively simple deep neural network based on the AlexNet architecture can be used to detect galactic rings from colour composite galaxy images. The training required only a modest sample size of 1122 original ringed galaxy images and about 10639 original images of galaxies without rings. To overcome the challenges of training with a small dataset of labelled rings, we used a smaller network and also increased the total number of images available for training through realistic image transformations, and during testing we used k-fold cross validation to make the best use of the data. We also used evaluation metrics suited for the class imbalance that was present in our data. We have prepared a catalog of candidate ring and non-ring galaxies using the predictions of our trained network. The probability score provided in the catalog can be used as a confidence measure of the presence of rings, to obtain smaller catalogues with the required level of confidence for follow up analysis. Additionally we used the network to generate a catalog of galaxies with rings as well a bars, which may be of special interest to galaxy morphologists.
Appendix A CASJOBS query
We provide the CASJOBS query used to create the catalog of ringed and non-ringed galaxies. The query is run in the DR18 context.
SELECT
  p.objid,p.ra,p.dec,
  (p.u-p.extinction_u) as umag,
  (p.g - p.extinction_g) as gmag,
  (p.r-p.extinction_r) as rmag,
  (p.i-p.extinction_i) as imag,
  (p.z - p.extinction_z) as zmag,
  (p.deVRad_g) as g_dev, p.deVRad_r, p.deVAB_g,
  s.specobjid, s.class,
  s.z as redshift into mydb.MyTable
  from PhotoObjAll AS p
JOIN SpecObj AS s ON s.bestobjid = p.objid
WHERE
  s.class=’GALAXY’
  AND s.z BETWEEN 0.01 and 0.2
  AND (p.g - extinction_g) < 16