### v2

# Automated Detection of Galactic Rings from SDSS Images

Abstract
Morphological features in galaxies, like spiral arms, bars, rings, tidal tails etc. carry information about their structure, origin and evolution. It is therefore important to catalog and study such features and to correlate them with other basic galaxy properties, the environment in which the galaxies are located and their interactions with other galaxies. The volume of present and future data on galaxies is so large that traditional methods, which involve expert astronomers identifying morphological features through visual inspection, are no longer sufficient. It is therefore necessary to use AI based techniques like machine learning and deep learning for finding morphological structures quickly and efficiently. We report in this study the application of deep learning for finding ring like structures in galaxy images from the Sloan Digital Sky Survey (SDSS) data release DR18. We use a catalog by Buta (2017) of ringed galaxies from the SDSS to train the network, reaching good accuracy and recall, and generate a catalog of 29420 galaxies of which 4855 have ring like structures with prediction confidence exceeding 90 percent. Using a catalog of barred galaxy images identified by Abraham et. al. (2018) using deep learning techniques, we identify a set of 2087 galaxies with bars as well as rings. The catalog should be very useful in understanding the origin of these important morphological structures. As an example of the usefulness of the catalog, we explore the environments and star formation characteristics of ring galaxies in our sample.
Astronomy data analysis (1858), Astronomy image processing (2306), Catalogs (205), Galaxies (573)
journal: ApJ
1 Introduction
Galaxies display a wide variety of forms as a consequence of the differences in their intrinsic structure, interaction with other galaxies as well as observational biases. Efforts to make sense of galaxy morphology with its multitude of forms have a long history dating back to Edwin Hubble and others. The Hubble classification system (Hubble, 1926) arranged galaxies in a sequence ranging from ellipticals to lenticulars to spirals to irregulars shaped objects. Galaxies other than the ellipticals were further classified into those with bar like structures and those without such structures, resulting in a tuning fork like diagram. Other more detailed classification systems have been developed (see, e.g., Van den Bergh, 1998), taking into account further morphological features.
In recent decades, the availability of many millions of higher resolution multiwavelength images from large deep field surveys with modern instrumentation have made possible detailed morphological studies (Buta, 2011). Galaxy morphology plays a key role in bettering our understanding of the secular processes that underlie galaxy evolution. Several important questions concerning the formation and evolution of galaxies can be addressed from the observed change in morphology with epoch.
A fairly common feature of disk galaxies is the presence of a ring shaped pattern in their light distribution (Buta & Combes, 1996). These rings in galaxies have come to be considered an integral part of galaxy morphology (Kormendy, 1979). Recognizing their importance, the de Vaucouleurs Revised Hubble-Sandage Classification System (VRHS; De Vaucouleurs, 1959) added rings as another dimension to the two-dimensional Hubble tuning fork, turning it into a classification volume. Rings in galaxies can fundamentally be divided into normal rings, also called ‘resonance rings’ and ‘catastrophic rings’, which are a result of galaxy collisions (Buta & Combes, 1996). Although different sub-classes exist within the catastrophic rings, such as polar, accretion, and collisional rings, they constitute a very small minority of all observed rings. Madore et al. (2009) have estimated the abundance of collisional rings to be only about 1 in 1000. Several theories have been put forward to explain the different ring structures that are observed, with the resonance interpretation being the most popular. The “Manifold Theory” that has been proposed Athanassoula et al. (2009) also has had considerable success in explaining several aspects of ring morphology.
Several catalogs that contain the visual morphology classification of galaxies are available today. Notable among them are Nair & Abraham (2010) and Fukugita et al. (2007), with the former containing the detailed classifications of 14,034 galaxies. Crowd-sourcing efforts such as the Galaxy Zoo citizen science project (Lintott et al., 2011) have managed to ramp up the number of galaxy morphology classifications to nearly 900,000 by involving volunteers. This vast database has motivated others to pick specific subtypes of galaxies for detailed classifications. For example, Buta (2017) has exhaustively classified 3962 ringed galaxies taken from this set. However, all these efforts involve a human in the loop for classifying galaxy morphologies.
With an explosion in the number of galaxy images being produced as part of imaging surveys like the Sloan Digital Sky Survey (SDSS; York et al., 2000), Panoramic Survey Telescope and Rapid Response System (Pan-STARRS; Kaiser et al., 2002), Dark Energy Survey (DES; Dark Energy Survey Collaboration et al., 2016) etc., automated methods of morphology classification have become very important. Although traditional image processing techniques, many of which are histogram-based, have had some success (see, e.g., Shamir, 2020), these are inappropriate for complex tasks such as shape recognition from images. However, the advent of artificial intelligence (AI) techniques like Convolutional Neural Networks (CNNs; LeCun et al., 1995) has significantly improved the automated efforts for galaxy morphology classification. One of the initial efforts in this area have been made by Banerji et al. (2010). Dieleman et al. (2015) have used CNNs to create a model that can match the combined consensus of the Galazy Zoo volunteers on all of the questions regarding galaxy morphology that are part of the survey. Abraham et al. (2018); Sánchez et al. (2018) have used CNNs for the automated classification of galaxy morphological features such as bars, bulges, edge-on morphology etc. More recently, Shimakawa et al. (2023) have used a deep-learning model to obtain ring classification for a very large sample of galaxies taken from the Hyper Suprime-Cam Subaru Strategic Program.
In the present study we investigate the efficacy of an automated method based on Convolutional Neural Networks for the detection of rings in galaxies. We use galaxy images from the Sloan Digital Sky Survey (SDSS; York et al., 2000) for the training and testing of our network. We treat the problem as a binary classification one, treating all sub-classes of rings as a single ring type and distinguishing it from galaxies that do not contain any rings, the non-ring type. The trained network can be used to identify ringed galaxies from the SDSS. We used the network’s predictions to generate a sample of 4855 ringed galaxies with a classification threshold of 0.90. Using this sample we explored the connection between ring formation and star formation activity considering the main sequence, green valley and quenched galaxy populations. We also analyzed the star formation by examining their distribution across different environments, categorized by their local surface densities.
The rest of the paper is organized as follows: In Section 2 we describe the data used in the analysis and introduce the neural network architecture in Section 3. We describe the data augmentation techniques and training procedure for the neural network in Sections 4 and 5 respectively and in Section 6 we analyse the results on the trained network. In Section 7 we present the catalog of ringed galaxies generated using our trained network and in Section 8 we present a subset of the ringed galaxies that have also been identified to have bars. In Section 9 we provide a discussion regarding the environments and star formation rates for the ringed galaxies in our catalog. Finally in Section 10 we give the concluding remarks.
2 Data
The performance of a supervised machine learning classifier depends to a large extent on the quality of the labelled data that is made available to it. For deep learning models the quantity becomes important as well because of the large number of trainable or free parameters in the model. Many of the standard datasets available for galaxy morphology classification do not contain rings as a class (see, e.g., DECALS10  [^1] https://astro.utoronto.ca/~hleung/shared/Galaxy10/Galaxy10_DECals.h5). Even with the Galaxy Zoo 2 dataset (Willett et al., 2013), ringed galaxies for which at least 50 percent of zoo volunteers agreed on a response are relatively small in number.
2.1 Rings and Non-Rings
We used the Buta (2017) catalog as our primary source to identify galaxies with rings. The author has provided detailed visual morphological classification of 3962 galaxies such galaxies. The classifications are done within the framework of the comprehensive de Vaucouleurs revised Hubble-Sandage system (CVRHS; Buta et al., 2015, 2007). The catalog also provides the author’s comments that are useful for creating a good quality training set by identifying rare cases of rings, poorly resolved galaxies etc. The galaxies used for classification by the author were picked from the Galaxy Zoo 2 (Willett et al., 2013) survey (hereafter, GZ2).
Along with a good quality training sample for rings, it is equally important to obtain a good sample of galaxies that do not have any presence of rings. Since the Buta (2017) catalog does not contain such galaxies, we used the Nair & Abraham (2010) catalog, which is one of the largest catalogs for the visual morphological classification of galaxies. They used galaxy images from the SDSS DR4 release (Stoughton et al. (2002); Adelman-McCarthy et al. (2006)). Galaxies with spectroscopic redshifts in the range $0.10.01<z<$ 0.1 and extinction corrected g’-band magnitude $<16$ were selected from the spectroscopic main sample (Strauss et al., 2002). This led to their final sample size of 14,034 galaxies.
Figure 1: The normalized distribution of redshifts and extinction corrected g-magnitude for the ringed galaxies(left) and non-ringed galaxies(right)
The classification scheme used in Nair & Abraham (2010) is not the CVRHS. Rather it is primarily based on the Carnegie Atlas of Galaxies (Sandage & Bedke, 1994) used in consultation with the Third Reference Catalog of Bright Galaxies (RC3; de Vaucouleurs et al., 1991). In order to use only galaxies without rings, we selected from the catalog galaxies that have the ring type and ring flag column set to zero.
2.2 Selection Criteria
The images used for the GZ2 survey come from the SDSS Data Release 7 (DR7; Abazajian et al., 2009). The GZ2 includes galaxies with extinction corrected Petrosian half-light magnitude in the r-band < $17.0$, Petrosian radius $\tt{petroR90_{r}}>3$ arcsec and spectroscopic redshift in the range $0.250.0005<z<$ 0.25, when it is known. Finally, galaxies that have SDSS flags which are either SATURATED, BRIGHT, OR BLENDED without an accompanying NODEBLEND were removed. We further selected galaxies that have an extinction corrected g’-band magnitude < 16 based on the selection criteria used for the non-ring galaxies. The spectroscopic redshift was also limited to be in the same range as that of the non-ring galaxies, specifically $0.10.01<z<$ 0.1. By visual inspection we removed 37 ringed and 797 non-ringed galaxies which contained artefacts and those which were not clearly distinguishable. After the selection criteria was applied and the manual removal, the number of ringed galaxies obtained were 1122 and the number of galaxies without rings came out to be 10,639. Figure 1 shows the distribution of redshift and extinction corrected g-magnitude for the ringed and non-ringed galaxies in our training set.
2.3 SDSS Image Cutouts
We modified a python library [^2] https://pypi.org/project/panstamps/ that exists in the public domain for downloading images we needed for training as well as catalog preparation. We incorporated multi-threaded parallel downloading using the Concurrent  [^3] https://docs.python.org/3/library/concurrent.futures.html python module into the script to speed up the downloading process. Figure 2 shows a sample of the galaxies with and without rings in our training set.
Galaxies with rings
Galaxies with no rings
Figure 2: A sample of galaxies with and without rings from our training set.
3 Network Design
AlexNet (Krizhevsky et al., 2012) is one of the earliest CNN architectures designed and yet one of the most reliable. Having a smaller size than most modern architectures results in lesser training time and also the requirement of lesser amount of training data for this network. The original AlexNet architecture was designed for the ImageNet challenge (Deng et al., 2009). The network has eight layers with the first five being convolutional layers and the remaining three being fully-connected. The first convolutional layer filters the 224 × 224 × 3 input image with 96 kernels of size 11 × 11 × 3 with a stride of four pixels. All of the convolution and fully connected layers are immediately followed by a block consisting of a Rectified Linear Unit activation (ReLU; Nair & Hinton 2010) and a batch normalisation (NORM) layer. The first, second and final convolution layers are each followed by a maxpooling layer, that leads to 50 percent downsampling of the convolution feature maps. These layers also contribute to the rotational invariance of the learnt model weights (Boureau et al., 2010). The max-pooling layers of the first two convolutions are followed by dropout layers (Srivastava et al., 2014) having a dropout fraction of 0.25. These prevent overfitting of the model by randomly dropping 25 percent of the weights during each training iteration.
We left the input layer size unchanged since it was suitable for our purposes. However, galactic rings are features that are harder to distinguish in comparison to the classes in the ImageNet challenge. We therefore used a smaller size of five for the kernels in the first convolutional layer. This seemed to have a considerable impact on the accuracy of the network. Thus in our network the first convolution layer has 96 kernels of size $5\times 5\times 3$ and the output is fed to the second layer of convolution with 256 kernels each of size $5\times 5\times 48$ The third, fourth and fifth convolutional layers have respectively 384, 384 and 256 kernels of size $3\times 3\times 256$, $3\times 3\times 192$ and $3\times 3\times 192$ respectively. These learn more complex features and their final output is fed into another max-pooling layer. The output of the final pooling layer is then fed into a series of fully connected layers of size 4096. These layers are connected by dropout layers which have 0.50 dropout fraction. The final layer is the sigmoid activation function layer which outputs a value that can be interpreted as the probability of the input belonging to the positive class. The architecture of our network is shown as a table and schematic diagram in Figure 3.
Name Type Shape
0 conv2d Conv2D (None, 100, 100, 96)
1 activation Activation (None, 100, 100, 96)
2 batch_normalization BatchNormalization (None, 100, 100, 96)
3 max_pooling2d MaxPooling2D (None, 49, 49, 96)
4 dropout Dropout (None, 49, 49, 96)
5 conv2d_1 Conv2D (None, 49, 49, 256)
6 activation_1 Activation (None, 49, 49, 256)
7 batch_normalization_1 BatchNormalization (None, 49, 49, 256)
8 max_pooling2d_1 MaxPooling2D (None, 24, 24, 256)
9 dropout_1 Dropout (None, 24, 24, 256)
10 conv2d_2 Conv2D (None, 24, 24, 384)
11 activation_2 Activation (None, 24, 24, 384)
12 batch_normalization_2 BatchNormalization (None, 24, 24, 384)
13 conv2d_3 Conv2D (None, 24, 24, 384)
14 activation_3 Activation (None, 24, 24, 384)
15 batch_normalization_3 BatchNormalization (None, 24, 24, 384)
16 conv2d_4 Conv2D (None, 24, 24, 256)
17 activation_4 Activation (None, 24, 24, 256)
18 batch_normalization_4 BatchNormalization (None, 24, 24, 256)
19 max_pooling2d_2 MaxPooling2D (None, 11, 11, 256)
20 dropout_2 Dropout (None, 11, 11, 256)
21 flatten Flatten (None, 30976)
22 dense Dense (None, 4096)
23 activation_5 Activation (None, 4096)
24 batch_normalization_5 BatchNormalization (None, 4096)
25 dropout_3 Dropout (None, 4096)
26 dense_1 Dense (None, 4096)
27 activation_6 Activation (None, 4096)
28 batch_normalization_6 BatchNormalization (None, 4096)
29 dropout_4 Dropout (None, 4096)
30 dense_2 Dense (None, 2)
31 activation_7 Activation (None, 2)
Figure 3: A tabular description of the AlexNet model architecture that we have used, generated with the Keras package (above) and a visual representation of the same (below). Code for generating the figure was adapated from Iqbal (2018)
3.1 Selection of Hyperparameters
Hyperparameters are those parameters that are not learnt or updated during the learning process and hence need to be manually set by the programmer. These can include the network architecture itself and several other parameters like the batch size used for the stochastic gradient descent or the number of epochs of training etc. Hyperparameter tuning is an important part of a deep learning project, since these parameters have a direct bearing on the performance and accuracy of the model.
The loss function, which specifies the error and how much weightage should be given to individual errors is an important hyperparameter to tune. We have used categorical crossentropy [^4] https://www.tensorflow.org/api_docs/python/tf/keras/metrics/categorical_crossentropy, which is a standard loss function suitable for binary classification problems. The optimizer function has an effect on how the network weights are updated in order to minimize this loss. We used the Adam optimizer with a learning rate of 0.003. Regularization techniques are a set of safeguard measures that prevent the model from overfitting on the training data and thus losing its ability to generalize. We have used L2 weight regularization with a factor of 0.0002 to penalize the weights from growing too much during training. The batch size controls how many input samples are considered together for evaluating the loss function. This leads to a trade-off between accuracy and speed. A batch size of 16 was seen to be sufficient in our case. The number of epochs or (num epochs) decides how many times the network gets to see the complete training data. We trained the network for num epochs without early stopping.
4 Image augmentation
Neural networks are prone to overfitting if the data available for training is insufficient. In our case the possibility of increasing the training data is limited by the availability of catalogs of ringed galaxies. It is therefore necessary to increase the training sample size, through data augmentation by transforming the available images. We have used operations such as horizontal and vertical flipping, rotation through arbitrary angles, brightness and contrast adjustments for augmenting our training set. Using rotations also makes the network invariant to rotations, which is needed because of the rotational symmetry of the ringed galaxies, The augmentations that we have used are illustrated in Figure 4 using a random galaxy.
We implemented image augmentations using the on-the-fly mode provided by the keras-tensorflow (Chollet et al., 2015) framework. Here a predefined sequence of transformations is applied to each image from the training set, using parameters which are randomized so that at each epoch, the network sees a different set of augmented images. In this way we avoid generating all the augmented images before the start of the training and also save disk space.
Original
Horizontal Flip
Vertical Flip
Rotation
Brightness
Contrast
Figure 4: The different types of image augmentations that are done to each image in our training set.
5 Training
The galaxy image cutouts used for training are 3 channel RGB $256\times 256$ pixel images which are are dynamically resized by the Keras image_dataset_from_directory  [^5] https://www.tensorflow.org/api_docs/python/tf/keras/utils/image_dataset_from_directory function into the required input shape of $240\times 240$. The pixel intensities in the range (0-255) are rescaled to the range (0,1) for easier computation. The input size and other parameters for the network like num epochs, batch size are controlled through a config file which is read at run time. The images are first resized to the desired input size and then loaded into memory in nbatches using the Keras image_dataset_from_directory function. A generator avoids the need for all of the images to be loaded into memory at once.
The data is initially split into a training and testing set using a ratio of 80:20. During training the training set is further split internally into a training and validation set using the same ratio. Before augmentations are applied, the number of ringed galaxies in our training set (including the validation split) is 897 and those in the testing set is 225. Similarly, the number of galaxies without rings in our training set is 8511 and those in the testing set is 2128. The network was trained for a total of 140 epochs. The training graph in Figure 5 shows the variation of the categorical_crossentropy loss function with increasing epochs for both the training and validation data. If there is overfitting, the validation loss of the network increases whereas the training loss keeps decreasing. In our case, the training loss and validation loss are seen to decrease together with each epoch and therefore there is no significant overfitting that occurs. There is further scope for training the network albeit with minimal returns.
Class Precision(%) Recall(%) F1-Score(%)
Rings 93 89 91
Non-Rings 99 99 99
Figure 5: (1) The training graph, which shows a plot of the training and validation loss against the number of epochs (top-left). (2) The Precision-Recall curve which shows the plot of precision against recall for ringed galaxies in our test set computed at various thresholds (top-right). The dashed line shows the curve expected from a no-skill classifier. A few thresholds taken at random are annotated on the curve. (3) The classification report showing the precision and recall for both rings and non-rings in our test set (bottom).
The training was done on the Amazon Web Services platform on an Intel Xeon workstation with 4 virtual CPUs and 16 GB RAM, making use of a Tesla (g4dn.xlarge) GPU accelerator with 16 GB graphics memory. A community AMI (Amazon Machine Image) running Arch Linux was modified to function as our base operating system. The data and model are version controlled using Data Version Control (DVC; Kuprieiev et al., 2022) and stored on Amazon S3 cloud storage. The code itself is version controlled and made available on github (Abraham, 2024). The entire training session takes 2.5 hours to run with our hardware specifications.
6 Results
6.1 Evaluation Metrics for Unbalanced Datasets
For evaluating the performance of the network in discriminating between galaxies with rings and those without rings, we us the three metrics accuracy, precision and recall. Accuracy, is the number of correct classifications as a fraction of the total number of classifications made.
$\rm{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$
where TP denotes the number of true positives, FP the number of false positives and FN the number of false negatives. Precision and recall are defined as
$\rm{Precision}=\frac{TP}{TP+FP}$
$\rm{Recall}=\frac{TP}{TP+FN}$
Precision is an indicator of the purity of the classification, that is the extent to which a collection of galaxies identified as ringed are indeed ring galaxies. Recall on the other hand is a measure of the extent of completeness in identifying galaxies from a sample containing galaxies with and without rings. Precision and recall are therefore important for astronomers.
Our trained model obtained a classification accuracy of 98 percent on the testing set. The threshold used for converting prediction scores to class labels is 0.50. The classification report shown in Figure 5 summarizes the precision and recall we have obtained for both the classes. The precision obtained for the rings was 93 percent. Our model also has a good recall value of 89 percent.Using the predicted probabilities of our network, a catalog with higher purity can be obtained by appropriately selecting the classification threshold. It is to be kept in mind that higher purity comes at a cost of reduced completeness. The nature of the trade-off can be understood using the precision-recall curve shown in Figure 5. Figure 6 shows galaxies in the test set that are correctly classified (True Positives and True Negatives) and failures (False Positives and False Negatives) obtained using our trained network.
True positives
True negatives
False positives
False negatives
Figure 6: Examples of images from our test set that are both correctly and incorrectly classified.
6.2 Prediction at Different Zoom Levels
To get a better insight into what the model has learnt we downloaded a single galaxy at different zoom levels using the SDSS image cutout service. This is done by varying the scale factor and downloading the different images. The trained network was then used to predict on these images. Figure 7 shows different versions of a particular ringed galaxy along with the probability score returned by the network. It can be seen that probability score goes to zero both when the images is too zoomed out or zoomed in such that the ring feature is not visible. For other zoom levels the probability score given by the network are comparable to what a human observer would assign.
Figure 7: Cutouts of a single galaxy downloaded at various zoom levels from the SDSS cutouts server and predicted using the trained network. The prediction scores obtained for each are shown in red color.
6.3 Analysis of Failed Predictions
To get a better idea of the performance of our classifier we analyzed the failed predictions. A quantitative way of doing this is to compute the log-loss for each sample in our test set. The log-loss is defined by
$\rm{logloss_{i}}=-[y_{i}\ln p_{i}+(1-y_{i})\ln(1-p_{i})]$
where $y_{i}$ is the true value which is either 0 or 1, and $p_{i}$ is the predicted probability score for observation $i$. In our binary classification problem, with a sigmoid activation for the last layer, the predicted value is a real valued number between 0 and 1, which is interpreted as a probability of the sample actually belonging to the positive class. For the negative class a low predicted probability implies a high confidence value. Hence in this case we need to take the difference of the value from one in order to get the real confidence. When analyzing the failures in our testing set we can use a loss function that is weighted using the confidence with which a prediction has been made.
The histogram of these losses can then be used to get an idea about the overall performance of the model. Figure 8 shows the histogram of log-loss obtained for our test set. The peak of the histograms are towards the lower end, which means that there are a lesser number of wrong predictions made with a high confidence value and greater number of correct predictions with a high confidence value. That is, it can be seen that the confidence of wrong prediction is much lower than that of the correct predictions. The figure also shows representative images from our test set corresponding to the false negatives and having high log-loss values. The rings in these galaxies are not very distinct and one of the galaxies has a high inclination angle that might have confused the network.
6.4 Model Interpretability
Once the model is trained, if we use it to make predictions on the training data itself, we will be able to get some idea of what the model has learnt, which is known as model interpretability. We used the trained model to predict the probability of each ringed galaxy in the sample being identified as ringed. The difference (1-predicted probability) can be taken as a measure of the error. This error can be used for finding subsets of ringed galaxies in our training sample that are either easy for the network to learn or difficult to learn. The results for a sample are shown in Figure 9. It can be seen that among the galaxy images that the network had trouble in learning, there are many with image artefacts or the presence of background stars or galaxies. Many of the galaxies have a zoomed in view as well. The galaxy images that were easily identified by the network had a better contrast with the background and the galaxies were at a normal zoom level.
Figure 8: The normalized distribution of log-loss in our test set (above). The true negatives are shown separately because of the difference in bin range. A sample of galaxies from the false negatives having the highest log-loss values (below). The annotations shown in red color denote the prediction scores.
Figure 9: Ringed galaxies from our training sample predicted using the trained model and with a low log-loss (above) and high log-loss (below). The annotations shown in red colour denote the prediction scores.
7 A Catalog of Ringed Galaxies
To prepare a catalog of ringed galaxies from SDSS we used the same selection criteria that was used for the training data preparation. The SDSS data can be queried using the CASJOBS  [^6] https://skyserver.sdss.org/casjobs/default.aspx interface. The complete query is given in the Appendix A. The source list obtained from the CASJOBS server was downloaded using our bulk download script. The trained network was then used to predict the ringed or non-ringed nature of these galaxies. A preview of the catalog is included in Table 1. Figure 10 shows a selection of galaxies with rings taken from our catalog.
Figure 10: Some good examples of ringed galaxies (above) and barred ring galaxies(below) from our catalog. The annotations shown in red colour denote the prediction scores.
In order to do a more thorough testing of our network on real world data, we show in Figure 11 an image mosaic created from 50 randomly selected galaxies from our prediction set along with the machine generated probability scores. The galaxies are sorted in descending order of the prediction score. Based on the natural abundance of ringed galaxies, we may expect less than 10 of these to be ringed. It can be seen on visual inspection that most of the galaxies with true rings have a high prediction scores in comparison to the non-rings and hence they are clustered towards the top of the image mosaic. We found 4855 galaxies that are predicted to have rings using a classification threshold of 0.90. This corresponds to a ring galaxy fraction of 16.5 percent. At a slightly lower threshold of 0.89 we recover the same ring galaxy fraction of 18.2 as observed by Nair & Abraham (2010). Figure 12 shows how this fraction varies for all thresholds between 0 and 1. Further in Figure 13 we show how the distribution of the extinction corrected g-magnitude of the predictions compares with that of the training sample.
Figure 11: A mosaic of 50 randomly selected images from the prediction set sorted using the predicted probability of having a ring. Visual inspection shows that the rings have accumulated towards the bottom and the non-rings towards the top.
ra dec objid gmag deVRad_r deVAB_g redshift Prediction Label
40.285690 -0.714957 1237645941824356443 15.641980 9.327478 0.339572 0.040287 0.908651 Rings
57.025337 0.208845 1237645942905438473 15.497970 10.815760 0.871453 0.025475 0.598131 NonRings
56.781387 1.000343 1237645943979114582 15.879010 2.804902 0.321389 0.039847 0.250911 NonRings
56.847420 0.875488 1237645943979114622 15.182350 13.953940 0.724238 0.039371 0.865321 NonRings
57.248385 0.925979 1237645943979311221 15.657840 6.806947 0.535690 0.035788 0.818097 NonRings
57.674720 1.040755 1237645943979507954 15.344130 3.541592 0.822157 0.037123 0.516564 NonRings
243.708876 -0.915653 1237648672921485632 15.293520 8.426456 0.640077 0.030767 0.836999 NonRings
246.015172 -0.902869 1237648672922468973 15.979040 15.339940 0.550030 0.046526 0.494671 NonRings
245.367353 -0.457074 1237648673459077169 15.901310 8.323227 0.500005 0.059092 0.534696 NonRings
246.782081 -0.492432 1237648673459667234 15.630960 8.579305 0.437323 0.046192 0.457613 NonRings
189.522249 -0.027031 1237648673971437623 14.639720 18.461840 0.442339 0.012513 0.802063 NonRings
237.945144 -0.105170 1237648673992671592 15.529860 5.215971 0.949471 0.054426 0.630027 NonRings
243.236782 -0.096251 1237648673994965546 14.331720 27.210810 0.723929 0.030867 0.879053 NonRings
243.583196 -0.031564 1237648673995162093 15.782320 5.114430 0.559130 0.030811 0.536166 NonRings
245.381888 -0.072364 1237648673995948107 15.426320 3.416426 0.888818 0.027287 0.904141 Rings
248.064172 -0.049932 1237648673997127724 15.503360 5.104637 0.811834 0.044017 0.924116 Rings
195.644361 0.348565 1237648674510995594 14.989240 7.022532 0.659088 0.067754 0.922850 Rings
242.672630 0.276520 1237648674531639653 15.742830 5.721440 0.605837 0.062059 0.449378 NonRings
243.154609 0.379972 1237648674531836029 15.808500 3.519223 0.956060 0.043145 0.945323 Rings
249.820694 0.410283 1237648674534719840 15.602230 4.794272 0.273508 0.024328 0.069281 NonRings
ra dec SDSS_Objid Prediction Label
$0$1.113000 1237663275780276407 0.001809 Barred-NonRings
0.019800 0.781700 1237657191978959126 0.650422 Barred-Rings
$0$0.723700 1237663783123681369 0.039325 Barred-NonRings
$0$1.213600 1237663275780276438 0.060768 Barred-NonRings
0.114500 14.962500 1237656495650570466 0.208050 Barred-NonRings
0.117000 14.381200 1237652942638481637 0.183115 Barred-NonRings
0.128700 -1.213000 1237663275780341888 0.532812 Barred-Rings
0.158900 14.623600 1237656495113699423 0.024679 Barred-NonRings
0.166500 -1.191200 1237663275780341936 0.036194 Barred-NonRings
0.228100 15.218700 1237652943712288939 0.000647 Barred-NonRings
0.241000 14.190400 1237656494576894171 0.162884 Barred-NonRings
0.241700 14.864400 1237652943175418025 0.215693 Barred-NonRings
0.261300 -0.257100 1237663783660617841 0.001189 Barred-NonRings
0.297800 15.985900 1237656496724443439 0.000518 Barred-NonRings
0.300100 15.351400 1237652943712288996 0.173043 Barred-NonRings
$0$0.992900 1237657189831606381 0.008202 Barred-NonRings
0.368800 -0.224400 1237663783660683434 0.321357 Barred-NonRings
0.380900 14.407500 1237652942638612625 0.211441 Barred-NonRings
0.388600 -0.730500 1237663783123812523 0.000156 Barred-NonRings
0.398900 -0.395200 1237663783660683462 0.303564 Barred-NonRings
Table 1: A preview of the catalog of ringed galaxies (above) and of barred rings (below) with the first 20 rows from each catalog shown.
Figure 12: Plot showing the fraction of ring galaxies predicted at various thresholds.
Figure 13: A comparison of the extinction corrected g-magnitude distribution between the training and the prediction samples shown for both the ringed galaxies (left) and galaxies without rings (right).
8 A Catalog of Ringed galaxies with Bars
Bars in galaxies are much more common than rings. Approximately half of the observed galaxies are seen to have bars (Knapen, 1999). However galaxies that contain both bars and rings are certainly a smaller fraction. Dedicated catalogs for such galaxies are also quite rare. Here we use the Abraham et al. (2018) catalog, which is an automated catalog of galaxies with bars in SDSS, to find galaxies with both bars and rings. The Abraham et al. (2018) catalog contains 111,838 galaxies that have been predicted as either barred or unbarred. We selected the barred galaxies which came to be 25,781 in number and used our trained network to predict which of these galaxies also have rings. Figure 10 shows a random selection of the barred galaxies that have been predicted to have rings with a probability score of greater than 90 percent. We found 2087 galaxies that contained both bars and rings using a classification threshold of $0.5$. A preview of the catalog containing barred galaxies with predictions for the presence or absence of a ring is shown in Table 1.
9 Discussion
As a demonstration of the usefulness of the catalogs generated in this paper, in this section, we will explore the environments and star formation characteristics of ring galaxies in our sample. Our goal is to analyze star formation within the framework of the main sequence of star-forming galaxies, the transitioning population of green valley galaxies, and the quenched galaxies. Additionally, we will examine star formation in relation to the environmental variations from low to high densities. For this analysis, we use a sample of 4855 galaxies with a classification threshold of 0.90 that is predicted to have rings.
[…]
Figure 14 presents the histogram of $\log\ \Sigma$ for the sample of 4471 ring galaxies with environment information. We find that ring galaxies are primarily found in intermediate density environments, i.e. in galaxy groups (49.3%) and in low-density environments, i.e. as isolated galaxies 35.1%). We see relatively few ring galaxies (15.6%) in high-density environments in agreement with Elmegreen et al. (1992). Numerical simulations, such as ROMULUS (Tremmel et al., 2017), suggest that galaxies within group environments often undergo interactions or multiple mergers with gas-rich satellites, evolving into elliptical galaxies. Subsequently, these galaxies start to accrete gas and redevelop their disks, which then become unstable, triggering star formation in ring-like patterns. When galaxies display rejuvenated star-forming rings, two pathways for gas inflow are identified in numerical simulations: (a) Diffuse cooling gas that feeds the galaxy, aiding in developing the disk/ring structure. (b)The acquisition of ram pressure-stripped gas streams from gas-rich satellites (Jung et al., 2022).
To study star formation in our sample ring galaxies with environment information, we utilized the GALEX-SDSS-WISE Legacy Catalog (GSWLC-X2) (Salim et al., 2016, 2018) and the cross-match gave us 4022 ring galaxies with a stellar mass (log M) range of 8.8-11.68. To investigate the influence of the environment on the star formation process in ring galaxies, we constructed a control sample of non-ring galaxies with a classification threshold of 0.40 for being a ring galaxy, i.e., a threshold of 0.60 for being a non-ring galaxy. This control sample has a similar r-band magnitude and redshift, and a cross-match with the GSWLC-X2 catalog gave $\sim$5000 non-ring galaxies. The GSWLC-X2 catalog integrates ultraviolet data from GALEX, optical data from SDSS, and infrared data from the WISE all-sky survey to model galaxy spectral energy distribution (SED). During SED modelling, the catalog applies flexible dust attenuation laws and emission line corrections to ensure accurate measurements of star formation rate (SFR) and stellar mass (M), as it might lead to the wrong identification of galaxies.
We calculated the specific star formation rate (sSFR) and, following Salim (2014), classified our sample galaxies into three sub-classes as a function of environment, as shown in Figure 15. The star-forming region is identified with log sSFR $\geq$ -10.8, the green valley, which is a transition zone between the star-forming and quenched regions, is defined as -11.8 < log sSFR < -10.8, and quenched region is defined by log sSFR $\leq$ -11.8. In high-density environments, most of the ring galaxies are in the quenched (42.3 %) or green valley region (33.6 %); however, the control sample of non-ring galaxies are almost equally distributed in quenched (33.8%) and star-forming region (47.8%) with fewer galaxies in green valley (18.9%). In contrast, low and intermediate-density environments are dominated by star-forming ring galaxies (52.4% and 43.2 %, respectively), as the non-ring galaxies from the control sample (86.9% and 76.3 %, respectively). However, the number of ring galaxies in green valleys increases for intermediate-density environments, though the number of non-ring galaxies increases for high-density environments. It is apparent that most of the ring galaxies are located in the star-forming region for low and intermediate-density environments, which validates the scenario in which rejuvenation of star formation happens by accreting gas and triggering star formation in ring-like features. Our results are somewhat different from those of Fernandez et al. (2024), who suggest that ringed galaxies exhibit lower star formation activity in intermediate-density environments, though it is much lower in high-density environments.
Many ring galaxies from our sample are also part of the MaNGA survey, and the study of spatially resolved stellar populations of rings using such IFU surveys will provide useful insight into the origin of star-forming rings, and this can be the subject of future work.
Figure 14: Histogram of local environmental density ($\log\ \Sigma\ (\text{Mpc}^{-2})$) for our final sample of galaxies. The dashed lines separate the sample into low ($\log\Sigma\ (\text{Mpc}^{-2})<-0.5$), intermediate(-0.5 < $\log\Sigma\ (\text{Mpc}^{-2})<0.5$), and high densities($\log\Sigma\ (\text{Mpc}^{-2})>0.5$).
Figure 15: Dependence of specific star formation rate on stellar mass (sSFR-Mass plane) for our sample of ring galaxies and control sample of non-ring galaxies (gray dot) as a function of (a) low-density ($\log\Sigma\ (\text{Mpc}^{-2})<-0.5$) (b) intermediate-density (-0.5 < $\log\Sigma\ (\text{Mpc}^{-2})<0.5$), and (c) high-density environment ($\log\Sigma\ (\text{Mpc}^{-2})>0.5$)
10 Conclusion
We have shown that a relatively simple deep neural network based on the AlexNet architecture can be used to detect galactic rings from colour composite galaxy images. The training required only a modest sample size of 1122 original ringed galaxy images and about 10639 original images of galaxies without rings. To overcome the challenges of training with a small dataset of labelled rings, we used a smaller network and also increased the total number of images available for training through realistic image transformations. We also used evaluation metrics suited for the class imbalance that was present in our data. We have prepared a catalog of candidate ring and non-ring galaxies using the predictions of our trained network. The probability score provided in the catalog can be used as a confidence measure of the presence of rings, to obtain smaller catalogs with the required level of confidence for follow up analysis. Additionally we used the network to generate a catalog of galaxies with rings as well a bars, which may be of special interest to galaxy morphologists. Finally, we explored the connection between the environments and star formation characteristics of the 4855 ringed galaxies in our catalog. We found that ring galaxies are star-forming and primarily found in low- and intermediate-density environments.
Appendix A CASJOBS query
We provide the CASJOBS query used to create the catalog of ringed and non-ringed galaxies. The query is run in the DR18 context.
SELECT
  p.objid,p.ra,p.dec,
  (p.u-p.extinction_u) as umag,
  (p.g - p.extinction_g) as gmag,
  (p.r-p.extinction_r) as rmag,
  (p.i-p.extinction_i) as imag,
  (p.z - p.extinction_z) as zmag,
  (p.deVRad_g) as g_dev, p.deVRad_r, p.deVAB_g,
  s.specobjid, s.class,
  s.z as redshift into mydb.MyTable
  from PhotoObjAll AS p
JOIN SpecObj AS s ON s.bestobjid = p.objid
WHERE
  s.class=’GALAXY’
  AND s.z BETWEEN 0.01 and 0.1
  AND (p.g - extinction_g) < 16